### v2

@@ 3.4 R4 Experiment
Inherent nondeterminism in ML models means that results can vary between test runs, even with identical code, data, and hyperparameters. This variation occurs from unanticipated influences during training, such as random parameter initialization, stochastic optimization, and random data subsampling (e.g., in k-fold cross-validation) [65, 44]. Neural networks are especially known for their inherent nondeterminism, leading to varied computational outcomes in multiple reruns due to numerous sources of randomness during training [1]. In some cases, inherent non-determinism can cause such large variations that reruns not only yield slightly different outcomes, but also lead to significant fluctuations in the performance of an ML model [1]. This issue is exacerbated when other sources of variation, such as different hyperparameters, are introduced to the ML model. In such cases, the impact can be magnified, and it is often observed that minor changes in hyperparameters can result in significant performance loss [11]. For example, a review of the status quo of reproducibility in NLP research highlights some core issues with non-determinism that are exemplary for ML research [11]. Simply rerunning the original code during a reproduction can lead to different computational outcomes compared to the originally reported ones. Reinforcement learning, a subfield of ML, is particularly susceptible to these reproducibility issues, partially because of additional sources of non-determinism in the learning process, such as the reinforcement learning environment or policy [52].
@@
Various studies have demonstrated that hardware differences, such as different GPUs or CPUs, and compiler settings can lead to different computational outcomes [34]. Additionally, a comparison between the same ML algorithm with fixed random seeds executed using PyTorch [^1] https://pytorch.org/ and TensorFlow [^2] https://www.tensorflow.org/ resulted in different performances [60]. A comparison of the results of an ML experiment between different ML platforms also shows that out-of-the-box reproducibility is not provided. In general, it is questionable to what extent reproducibility can be ensured out-of-the-box [30]. Another important factor is the use of GPUs, which can increase randomness compared to the use of CPUs. This is due to parallel optimization and the use of optimizers in ML frameworks, such as PyTorch and TensorFlow. As a result, some researchers have resorted to solely using CPUs for executing their experiments. However, this comes at the expense of runtime-efficiency [3].
@@
Another barrier to ML reproducibility is the limited access to computational resources, which most recently has become evident in the case of transformer-based large language models (LLMs) [10, 82]. These transformer architectures need a vast amount of data and computational resources, to which most researchers have limited access. Estimates have calculated the costs to reproduce one model to be around \$1 million to \$3.2 million, which largely increases the difficulty of reproducing results [10]. Another study found that the needed computational resources are one of the most significant factors impacting reproducibility [64]. Especially ML models, which require computational clusters for training and optimization, are notably hard to reproduce.
@@ 4 Drivers for ML Reproducibility
In this section, we discuss drivers for ML reproducibility, which we subdivide into (i) technology-based drivers, (ii) procedural drivers, and (iii) drivers related to awareness and education.
@@ 4.1 Technology-based Drivers
Utilizing hosting services offers an efficient way to share code, data, and ML model parameter settings, thus supporting the reproducibility of ML-driven research [74]. Examples of hosting services include the runtime environments of ML platforms. If the original author runs the ML experiment in such a runtime environment, e.g., Kaggle Notebooks [^3] https://www.kaggle.com/code, Google Colab [^4] https://colab.google/, or CodaLab [^5] https://codalab.org/, researchers attempting to reproduce the results should be able to execute the experiment within the same environment.
The main advantage of using a hosting service is that the provider takes care of the logistics of code hosting and distribution. However, the main drawbacks are the limits on data size and computational resources. Since these hosting services are run in the cloud, there are restrictions on how many resources a single user can utilize. The limit on resources varies between different hosting services and is sometimes tied to subscription levels. Because of these limits, hosting services may not be suitable for all research purposes, especially considering the compute-intensive nature of state-of-the-art ML models, such as LLMs. Furthermore, research has indicated that out-of-the-box reproducibility across multiple ML platforms is not yet guaranteed [30].
@@
Reproducing the environment and setup of any ML experiment requires the consideration of existing dependencies and software versions, and is usually a complex task itself. Virtualization can simplify this process by bundling the essential components of ML models and experiments, such as the environment and code, into a single package for being shared with other researchers. Thus, if the authors of a paper build the experiment in a virtual environment, issues associated with setup reproduction can be greatly reduced. However, the adoption of virtualization by researchers depends on its user-friendliness and the effort of integration into their current workflows [12]. Concerns about virtualization include its limitations in allowing researchers to build upon them in a scalable manner. Traditional virtual machines (VMs) emulate an entire operating system for setting up and running experiments. The use of containerization software like Docker [^6] https://www.docker.com/ has just recently become popular. Containers are more lightweight and flexible than VMs, making it easier to adapt environments for follow-up studies [12]. There are also designated platforms for computational research, such as Code Ocean [^7] https://codeocean.com/, that offer virtualization via so-called reproducible capsules. Their focus in particular is to simplify the virtualization process and allow researchers to focus on the research questions instead of the standardization of environments [16]. Additionally, there are many other tools, such as e.g., ReproZip [15], which is one of the recommended tools by the SIGMOD Reproducibility [^8] https://reproducibility.sigmod.org/ to streamline reproducibility, and DetTrace [53], which aims to ensure completely deterministic computations.
@@
Many different sources of randomness during ML training lead to irreproducibility of ML research. Managing these sources of randomness, e.g., via random number seeds, deterministic algorithms, or other methods, could therefore greatly increase reproducibility. Fixed andom number seeds should be used and published to make ML experiments more reproducible and control a number of sources of the inherent nondeterminism. A seed is a first value used to initialize the pseudo-random number generator. When the same seed is used, the sequence of pseudo-random numbers generated is deterministic, meaning it will be the same every time the code is run. Additionally, ML models should be benchmarked and evaluated with multiple random number seeds, such that the variance can be reported and inform about the true performance of an ML model [65]. However, fixing the random number seeds can only eliminate algorithmic sources of randomness. Furthermore, hardware nondeterminism is also avoidable by using patching, i.e., replacing non-deterministic operations from hardware with deterministic ones, however, this will lead to a significant computing overhead [14]. To counteract inherent nondeterminism in reinforcement learning and achieve reproducible evaluations, there are frameworks, e.g. Gym-Ignition [22, 13], rl_reach [8], or MinAtar [81], which act as standardized benchmarking environments. Within them, different algorithms designed for the same task can be evaluated and compared against each other in a common environment. The frameworks counteract the effects of inherent nondeterminism by automatically controlling random seeds and evaluating algorithms over a number of runs. Additionally, it is an ongoing field of research to implement fully deterministic reinforcement learning algorithms [52] and make use of them within such frameworks [73].
@@
Privacy-preserving technologies support reproducibility, as they enable the collaborative training of ML models without sharing private or sensitive data. The main benefit of this is, that ML models can make use of larger, and more diverse data, thus helping to decrease bias and leading to more reproducible ML models. The main aim of Privacy-Preserving Machine Learning (PPML) is to facilitate the use of privacy-sensitive data to create better ML models, and, to allow data owners to collaboratively train ML models on private data. In that regard, PPML has several requirements. First, protecting the confidentiality of the training data. Second, preventing the leakage of sensitive information from ML model parameters and outputs, i.e., to hinder the re-identification of individuals. Third, achieving the listed security and privacy aspects while still preserving the utility of the ML model [79]. To achieve this, a number of different techniques are being used and developed, such as differential privacy, fully homomorphic encryption, secure multi-party computation, and federated learning [20]. These techniques are implemented in software libraries such as TensorFlow Privacy, PySyft, ML Privacy Meter, CryptFlow, or Crypten [7]. Furthermore, data can be made anonymous by removing identifiable personal information. However, if too much data is removed, the ML models may perform poorly. If not enough data is removed, it may still be possible to re-identify individuals by combining many different non-unique features [79]. An alternative approach to PPML is to generate synthetic data that captures the same information as the original data. A robust technique for creating such datasets can produce readily available datasets of nearly any size, as demonstrated in Biomedical fields. This approach has led to the development of Synthea, a software package designed to generate synthetic patient data and electronic health care records [76].
@@
  • DVC [^9] https://dvc.org/ is a version control system for ML projects with a command-line interface similar to Git [^10] https://git-scm.com/. It integrates with Git, supports cloud storage, and handles large versioning of datasets. DVC ensures full code and data provenance by enabling experiment tracking.
  • MLflow [^11] https://mlflow.org/ is an open-source tool for supporting ML experiment tracking, ML model deployment, and centralized model storage. Additionally, it provides an easy-to-use Web dashboard.
  • RO-Crate is a specification, implemented by a number of tools, aimed at aggregating and describing research data and metadata [71]. Although not specifically designed for ML, RO-Crate can aggregate and represent any resource, making it applicable for managing ML artifacts as well.
  • dToolAI [32] collects and packages ML models together with supplemental information, such as hyperparameter settings, appropriate metadata, and persistent URIs for model training data. In contrast to the other tools, dToolAI is specifically tailored for deep learning models.
AutoML platforms, such as H2O Driverless AI [^12] https://h2o.ai, Google Cloud AutoML [^13] https://cloud.google.com/automl, DataRobot [^14] https://www.datarobot.com/platform) are a novel subcategory of ML tools that aim to aid with every aspect of the ML lifecycle, from data aggregation to model deployment. Thus, AutoML tools could facilitate more standardized ML models and also take care of tasks like hyperparameter optimization. It is however questionable how practical these tools are for reproducible ML research, since they often hide ML model optimization procedures. Sometimes, e.g., in the case of H2O Driverless AI, the features go a step further and promise to be able to prevent problems such as overfitting or target leakage [^15] https://h2o.ai/wiki/target-leakage//. In the case of target leakage, this is done by checking for a strong correlation between a feature and the target and then taking action, e.g., warning the user or automatically handling it. This is, however, a very simple solution to the problem and does not address the more complex cases of data leakage that are often present in research, e.g., temporal leakage.
@@ 4.2 Procedural Drivers
Due to a lack of shared datasets, currently, many researchers in ML-driven research - most notably in Biomedical fields - use individually acquired data [46]. The collection of such data is a time-consuming task and bears a significant risk of causing reproducibility issues, e.g., bias or data leakage. Often, the number of participants within the datasets is not very large and, thus, might not translate well to new contexts. Creating shared and standardized datasets can, therefore, (i) save researchers time in acquiring new data, (ii) facilitate the collaborative and independent maintenance and verification of data to minimize methodological errors, and (iii) support transferability and generalizability through the use of multi-institutional data [46]. In addition to the standardization of datasets, data cards [62] provide a consistent and comparable framework for reporting essential aspects of ML datasets. This includes information, e.g., about access restrictions, risks and limitations associated with the usage of the dataset, or any preprocessing steps, amongst many other contents, which are needed for reproducible ML development.
@@
Another issue is the lack of standardized evaluation methods, which leads to reported performances of ML models often being overly optimistic [19]. To ensure the statistical significance of ML model evaluations, it is crucial to report performance as an aggregate of results obtained from multiple random runs, and with different random number seeds [17]. Furthermore, ML models should, if possible, be tested and evaluated on multiple different datasets [21]. This underscores the need for standardized evaluation methods, which can be supported by checklists or tools to prevent errors in this critical aspect of ML research. For this, similarly to data cards, model cards [48] are aimed to standardize the evaluation and reporting of the performance of ML models for a variety of use cases. Model cards should inform users about the possible applications of the ML model and its limitations. For the creation of the model cards, a toolkit is also being developed by Google [^16] https://research.google/blog/introducing-the-model-card-toolkit-for-easier-model-transparency-reporting/. In reinforcement learning, the creation of standardized evaluation pipelines is continually being researched to enable reproducible benchmarking of different reinforcement learning algorithms [41].
@@
There are many guidelines and checklists that outline best practices for increasing the reproducibility of ML. The guidelines are often aimed at specific parts of the ML workflow. For example, the FAIR guiding principles [^17] https://www.go-fair.org/fair-principles/ aim to improve the management and stewardship of scientific data by making scientific data (i) findable, (ii) accessible, (iii) interoperable, and (iv) reusable. Other guidelines promote the transparency and openness of scientific reporting in general, such as the TOP guidelines [^18] https://www.cos.io/initiatives/top-guidelines, which target journals. Similarly, checklists provide a simple framework for ensuring certain criteria are met. Checklists have been applied effectively in the past, e.g., in safety-critical systems, where they were used as early as in 1935 to complete pre-flight checks in Boeing airplanes. A promising example is the ML checklist proposed in [59], which has been suggested as best practice by researchers of different fields, e.g., in chemistry [6]. The checklist requests information about (i) the models and algorithms being used, (ii) theoretical claims in the research article, (iii) data, (iv) code, and (v) the ML experiment(s). However, one drawback of reproducibility checklists when used for academic conferences and journals is the additional workload they impose on reviewers. To mitigate this workload, one option would be to leverage LLMs to assist the review process [45]. Finally, numerous guidelines and checklists for ML reproducibility have been recommended in various research fields [6]. Especially in Biomedical fields, we have seen a considerable adoption of guidelines and checklists, such as the TRIPOD statement [18], the CLAIM checklist [49], the ROBUST-ML checklist [2], or PROBAST [78].
4.2.3 Model info sheets and model cards.
@@
Model cards [48] are documentation sheets that provide information about ML models, including their intended use, potential limitations, and ethical considerations. They aim to enhance transparency in AI, by detailing aspects such as data used for training, performance metrics, evaluation methodologies, and possible biases. Model cards help users to understand and evaluate ML models more comprehensively, such that they are not being deployed in contexts other than the intended use, and thus to increase reproducibility. Similarly, model info sheets also provide documentation about ML models, but are specifically designed for the detection and prevention of data leakage in ML models [40]. The model info sheets are published alongside research to enable other researchers to quickly verify the validity of the data used to train ML models. Model info sheets require authors to answer detailed questions about the data and corresponding train/test splits, targeting various types of data leakage [40]. Despite that, they have two main drawbacks: first, verifying the correctness of info sheets only works after reproducing the results; second, completing these sheets requires a certain level of expertise in ML. In general, model cards and model info sheets represent a promising, low-effort driver for ML reproducibility. They are especially useful in handling some of the methodological issues associated with ML models that could arise, e.g., when being developed by non-experts [40].

### v3

@@ 3.4 R4 Experiment
Inherent nondeterminism in ML models means that results can vary between test runs, even with identical code, data, and hyperparameters. This variation arises from sources of randomness during training, such as, e.g., random parameter initialization, stochastic optimization, and random data subsampling (e.g., in k-fold cross-validation) and the complex interactions between them, which are fundamental characteristics of most ML workflows [101, 74]. Neural networks are especially known for their inherent nondeterminism, leading to varied computational outcomes in multiple reruns due to increased sources of randomness during training [1]. In some cases, inherent nondeterminism can cause such large variations that reruns not only yield slightly different outcomes, but also lead to significant fluctuations in the performance of an ML model [1] or to varying conclusions in ML model comparisons [47]. This issue is exacerbated when other sources of variation, such as different hyperparameters, are introduced to the ML model. In such cases, the impact can be magnified, and it is often observed that minor changes in hyperparameters can result in significant performance loss [14]. Reviews of reproducibility in both NLP research [14] and biomedical research [2] highlight these core issues with nondeterminism that are exemplary for ML research because they reflect challenges that are prevalent across ML-based research and serve as representative examples of the broader reproducibility crisis faced by the ML community. Simply rerunning the original code of an experiment during a reproduction leads to large variances of results and different computational outcomes on each run. Reinforcement learning, a subfield of ML, is particularly susceptible to these reproducibility issues, partially because of additional sources of nondeterminism, such as the reinforcement learning environment or policy [85].
@@
Various studies have demonstrated that hardware differences, such as different GPUs or CPUs, and compiler settings can lead to different computational outcomes [54]. Additionally, a comparison between the same ML algorithm with fixed random seeds executed using PyTorch [^5] https://pytorch.org/ and TensorFlow [^6] https://www.tensorflow.org/ resulted in different performances [96]. Furthermore, even different versions of the same framework can lead to different performance results [109]. A comparison of the results of experiments performed on different hosted ML platforms also found that out-of-the-box reproducibility is not guaranteed there [46]. Another important factor is the use of GPUs, which can increase randomness compared to the use of CPUs. This is due to parallel optimization and the use of optimizers in ML frameworks, such as PyTorch and TensorFlow. As a result, some researchers have resorted to solely using CPUs for executing their experiments. However, this comes at the expense of runtime-efficiency [4].
@@
The barrier to ML reproducibility posed by limited access to computational resources has recently become evident in the case of transformer-based Large Language Models (LLMs) [13]. These transformer architectures need a vast amount of data and computational resources, to which most researchers have limited access. Estimates have calculated the costs to reproduce one model to be around \$1 million to \$3.2 million [13]. Another study found that the needed computational resources are one of the most significant factors impacting reproducibility [100]. Especially ML models, which require computational clusters for training and optimization, are notably hard to reproduce.
@@ 4 Drivers for ML Reproducibility
In this section, we discuss drivers for ML reproducibility, which we subdivide into (i) technology-based drivers, (ii) procedural drivers, and (iii) drivers related to awareness and education. For every driver, we also provide case studies or examples from the literature illustrating the effectiveness of the driver for ML reproducibility.
@@ 4.1 Technology-based Drivers
Utilizing hosting services offers an efficient way to share code, data, and ML model parameter settings, thus supporting the reproducibility of ML-driven research [114]. Examples of hosting services include the runtime environments of ML platforms. If the original author runs the ML experiment in such a runtime environment, e.g., Kaggle Notebooks [^7] https://www.kaggle.com/code, Google Colab [^8] https://colab.google/, or CodaLab [^9] https://codalab.org/, researchers attempting to reproduce the results should be able to execute the experiment within the same environment. The main advantage of using a hosting service is that the provider takes care of the logistics of code hosting and distribution. However, the main drawbacks are the limits on data size and computational resources. Since these hosting services are run in the cloud, there are restrictions on how many resources a single user can utilize. The limit on resources varies between different hosting services and is limited by users’ available funds and sometimes subscription levels. Because of these limits, hosting services may not be suitable for all research purposes, especially considering the compute-intensive nature of novel ML models, such as LLMs.
As said above, the degree to which such services offer out-of-the-box reproducibility remains highly questionable [46]. Nevertheless, some hosting services have effectively been used to create end-to-end reproducible AI pipelines, especially in conjunction with standardized datasets such as the National Cancer Institute Imaging Data Commons [37]. The effectiveness of hosting services for ML reproducibility has been shown for research in radiology [18] and similar reproducibility experiments have been successfully conducted in pathology research [106]. As also noted in these experiments, it is important that results are reported with a quantification of the variance across different test runs, since effects of randomness could still be prevalent.
@@
Reproducing the environment and setup of any ML experiment requires the consideration of existing dependencies and software versions, and is usually a complex task itself. Virtualization can simplify this process by bundling the essential components of ML models and experiments, such as the dependencies and code, into a single package for sharing with other researchers. Thus, if the authors of a paper build the experiment in a virtual environment, issues associated with setup reproduction can be greatly reduced. However, the adoption of virtualization by researchers depends on its user-friendliness and the effort of integration into their current workflows [17]. Concerns about virtualization include its limitations in allowing researchers to build upon them in a scalable manner. Traditional virtual machines (VMs) emulate an entire operating system for setting up and running experiments. The use of containerization software like Docker [^10] https://www.docker.com/ has become more popular in recent years. Containers are more lightweight and flexible than VMs, making it easier to adapt environments for follow-up studies [17]. There are also designated platforms for computational research, such as Code Ocean [^11] https://codeocean.com/, that offer virtualization via so-called reproducible capsules. Their focus, in particular, is to simplify the virtualization process and allow researchers to focus on the research itself rather than the standardization of environments [25]. Additionally, there are many other tools, such as ReproZip [23], one of the recommended tools by the SIGMOD Reproducibility Availability and Reproducibility Initiative [^12] https://reproducibility.sigmod.org/ to streamline reproducibility, and DetTrace [86], which aims to ensure completely deterministic computations.
The use of containers is rapidly gaining in popularity across many research fields, e.g., neuroscience and genomics [82]. Notably, the platform Code Ocean has been integrated into the peer reviewing process by Nature journals to support the submission process of experiments [34]. This widespread adoption highlights the suitability of containers to enhance experiment sharing and improve reproducibility. Furthermore, a case study has compared ten different containerization-based approaches for reproducibility [24]. The strengths and weaknesses of each approach were analyzed, with results demonstrating the suitability of containers to enhance reproducibility by encapsulating the computational environment and to decrease the effort for publishing reproducible ML-based experiments.
@@
Many different sources of randomness during ML training lead to the irreproducibility of ML research. Managing these sources of randomness, e.g., via random number seeds, deterministic algorithms, or other methods, could therefore greatly increase reproducibility. Fixed random number seeds should be used and published to make ML experiments more reproducible and control a number of sources of inherent nondeterminism. A seed is a first value used to initialize the pseudo-random number generator. When the same seed is used, the sequence of pseudo-random numbers generated is deterministic, meaning it will be the same every time the code is run. Experiments have shown that fixing random seeds can effectively ensure reproducible results when algorithms are not being executed in parallel on GPUs [1]. Additionally, one case study has shown that achieving reproducibility for GPU-trained neural networks [21] is possible through a method known as patching. Patching aims to replace non-deterministic operations with deterministic ones. In the case study, a systematic patching approach was successful in achieving reproducible image classification results for six different neural network. However, this process also leads to higher computational costs and a time overhead, which was also analyzed in the study. Additionally, ML models should be benchmarked and evaluated with multiple random number seeds, such that the variance can be reported and inform about the true performance of an ML model [101]. Similarly, the use of uncertainty-aware quantification metrics to evaluate ML models can also help increase reproducibility [96].
Additionally, to counteract inherent nondeterminism in reinforcement learning and achieve reproducible evaluations, there exist frameworks, such as Gym-Ignition [38, 19], rl_reach [9], or MinAtar [121], which act as standardized benchmarking environments. Within them, different algorithms designed for the same task can be evaluated and compared against each other in a common environment. Some frameworks counteract the effects of inherent nondeterminism by automatically controlling random seeds and evaluating algorithms over a number of runs. Finally, it is an ongoing field of research to implement fully deterministic reinforcement learning algorithms [85] and make use of them within such frameworks [113].
@@
Privacy-preserving technologies support reproducibility, as they enable the collaborative training of ML models without sharing private or sensitive data. The main benefit of this is, that ML models can make use of larger and more diverse data, thus helping to decrease bias and leading to more reproducible ML models. The main aim of Privacy-Preserving Machine Learning (PPML) is to facilitate the use of privacy-sensitive data to create better ML models, and, to allow data owners to collaboratively train ML models on private data. In that regard, PPML has several requirements. First, protecting the confidentiality of the training data. Second, preventing the leakage of sensitive information from ML model parameters and outputs, i.e., to hinder the re-identification of individuals. Third, achieving the listed security and privacy aspects while still preserving the utility of the ML model [119]. To achieve this, a number of different techniques are being used and developed, mainly Differential Privacy (DP), Homomorphic Encryption (HE), Secure Multi-Party Computation (SMPC), and Federated Learning (FL) [30]. These techniques are implemented in software libraries such as TensorFlow Privacy, PySyft, ML Privacy Meter, CryptFlow, or Crypten [8]. Furthermore, data can be made anonymous by removing identifiable personal information. However, if too much data is removed, the ML models may perform poorly. If not enough data is removed, it may still be possible to re-identify individuals by combining many different non-unique features [119].
An alternative approach to PPML is to generate synthetic data that captures the same information as the original data. A robust technique for creating such datasets can produce readily available datasets of nearly any size, as demonstrated in biomedical fields. This approach has led to the development of Synthea, a software package designed to generate synthetic patient data and electronic health care records [116]. It is, however, important to mention that there is still a gap in efficiency between theoretical advancements and real-world applications when using PPML techniques. To this end, one study conducting reproductions of 26 state-of-the-art applications of PPML has highlighted the challenges of balancing computational efficiency, privacy guarantees, and model utility, while emphasizing the need for improved reproducibility, open-source availability, and practical scalability [65].
@@
  • DVC [^13] https://dvc.org/: A version control system for ML projects with a command-line interface similar to Git [^14] https://git-scm.com/. It integrates with Git, supports cloud storage, and handles large versioning of datasets. DVC ensures full code and data provenance by enabling experiment tracking.
  • MLflow [^15] https://mlflow.org/: An open-source tool for supporting ML experiment tracking, ML model deployment, and centralized model storage. Additionally, it provides an easy-to-use Web dashboard.
  • RO-Crate [^16] https://www.researchobject.org/ro-crate/: A specification, implemented by a number of tools, aimed at aggregating and describing research data and metadata [111]. Although not specifically designed for ML, RO-Crate can aggregate and represent any resource, making it applicable for managing ML artifacts as well.
  • dToolAI [52]: Collects and packages ML models together with supplemental information, such as hyperparameter settings, appropriate metadata, and persistent URIs for model training data. In contrast to the other tools, dToolAI is specifically tailored towards Deep Learning models.
AutoML platforms, such as H2O Driverless AI [^17] https://h2o.ai, Google Cloud AutoML [^18] https://cloud.google.com/automl, DataRobot [^19] https://www.datarobot.com/platform) are a novel subcategory of ML tools that aim to aid with every aspect of the ML lifecycle, from data aggregation to model deployment. Thus, AutoML tools could facilitate more standardized ML models and also take care of tasks like hyperparameter optimization. It is, however, questionable how practical these tools are for reproducible ML research, since they often hide ML model optimization procedures. Recent assessments of the reproducibility of AutoML tools also came to the conclusion that current platforms cannot provide out-of-the-box-reproducibility [46, 94]. In a qualitative analysis of the reproduction experiments, the latter study did identify areas in which such tools can be enablers for reproducibility, e.g., due to their automatic documentation capabilities. However, the authors also identified aspects that need to be addressed, such as the need for simplified tool user interfaces - as many participants were overwhelmed by tool complexity and could not make use of the documentation capabilities - and more built-in reproducibility capabilities, which support the sharing of code and data [94]. Furthermore, some AutoML tools, such as H2O Driverless AI, aim to address problems such as model overfitting. In the case of data leakage, this is done by checking for a strong correlation between a feature and the target and then taking action, e.g., warning the user or automatically handling it. This is, however, a very simple solution to the problem and does not address the more complex cases of data leakage that are often present in research, e.g., temporal leakage.
@@ 4.2 Procedural Drivers
Due to a lack of shared datasets, many researchers in ML-driven research - most notably in biomedical fields - have to use individually acquired data [78]. The collection of such data is a time-consuming task and bears a significant risk of causing reproducibility issues, e.g., bias or data leakage. Often, the number of individual participants represented within datasets is not very large and, thus, findings might suffer poor generalizability. Creating shared and standardized datasets can, therefore, (i) save researchers time in acquiring new data, (ii) facilitate the collaborative and independent maintenance and verification of data to minimize methodological errors, and (iii) support transferability and generalizability through the use of multi-institutional data [78]. In addition to the standardization of datasets, data cards [98] provide a consistent and comparable framework for reporting essential aspects of ML datasets. This includes information, e.g., about access restrictions, risks and limitations associated with the usage of the dataset, or any preprocessing steps, amongst many other contents, which are needed for reproducible ML development.
@@
Another issue is the lack of standardized evaluation methods, which leads to reported performances of ML models often being overly optimistic [29]. To ensure the statistical significance of ML model evaluations, it is crucial to report performance as an aggregate of results obtained from multiple random runs, and with different random number seeds [26]. Furthermore, ML models should, if possible, be tested and evaluated on multiple different datasets [31]. This underscores the need for standardized evaluation methods, which can be supported by checklists or tools to prevent errors in this critical aspect of ML research. For this, similarly to data cards, model cards [80] are aimed to standardize the evaluation and reporting of the performance of ML models for a variety of use cases. Model cards should inform users about the possible applications of the ML model and its limitations. In 2020, Google introduced the Model Card Toolkit for the creation of model cards [^20] https://research.google/blog/introducing-the-model-card-toolkit-for-easier-model-transparency-reporting/. In reinforcement learning, the creation of standardized evaluation pipelines is continually being researched to enable reproducible benchmarking of different reinforcement learning algorithms [66].
The National Cancer Institute Imaging Data Commons is an established example of standardized datasets in biomedical research [37]. As a cloud-based repository, it contains a collection of cancer imaging data and has been used successfully in reproduction experiments in combination with hosting services. Other notable examples include the MIMIC [62] database for electronic health records, or OGB [56] for applying ML on graph data. Efforts are also being invested into increasing the reproducibility of language models, e.g., with the Holistic Evaluation of Language Models (HELM) [75], which offers a broad, scenario-diverse, and multi-metric benchmarking suite for language models. As demonstrated by the authors, using HELM, new language models can be evaluated in a more comprehensive way. Furthermore, the Language Model Evaluation Harness (lm-eval) toolkit [16] is a framework designed for language model evaluations and concerned with reproducibility aspects. This tool has already been used by other researchers for more reproducible language model evaluations [36, 71]. However, it is important to recognize how irreproducible most major models currently are. There exists a live-tracker of model openness [^21] https://opening-up-chatgpt.github.io/, which has reported that many projects, even those claiming to be open source, “inherit undocumented data of dubious legality”, that few projects share data or model or human reinforcement learning (RLHF) weights, and that “careful scientific documentation is exceedingly rare” [76].
@@
There are many guidelines and checklists that outline best practices for increasing the reproducibility of ML. The guidelines are often aimed at specific parts of the ML workflow. For example, the FAIR principles [^22] https://www.go-fair.org/fair-principles/ aim to improve the management and stewardship of scientific data by making scientific data findable, accessible, interoperable, and reusable. Other guidelines promote the transparency and openness of scientific reporting in general, such as the TOP guidelines [^23] https://www.cos.io/initiatives/top-guidelines, which target journals. Similarly, checklists provide a simple framework for ensuring certain criteria are met. Checklists have been applied effectively in the past, e.g., in safety-critical systems, where they were used as early as in 1935 to complete pre-flight checks in Boeing airplanes. A promising example is the ML checklist proposed in [93], which has been suggested as best practice by researchers of different fields, e.g., in chemistry [7]. The checklist requests information about (i) the models and algorithms being used, (ii) theoretical claims in the research article, (iii) data, (iv) code, and (v) the ML experiment(s). However, one drawback of reproducibility checklists when used for academic conferences and journals is the additional workload they impose on already overburdened reviewers. To mitigate this, one suggestion is to leverage LLMs to assist the review process [77].
Finally, numerous guidelines and checklists for ML reproducibility have been recommended in various research fields [7]. Especially in biomedical fields, there has been a considerable adoption of guidelines and checklists, such as the TRIPOD statement [28], the CLAIM checklist [81], the ROBUST-ML checklist [3], or PROBAST [118]. A systematic review in biomedical research has shown that the use of checklists is linked to increased reporting quality [49]. The review examined 943 articles over two years and found that mandatory checklists increased the inclusion of the main methodological information needed to reproduce the experiments by 65%.
4.2.3 Model cards and model info sheets.
@@
Model cards [80] are documentation sheets that provide information about ML models, including their intended use, potential limitations, and ethical considerations. They aim to enhance transparency in AI, by detailing aspects such as data used for training, performance metrics, evaluation methodologies, and possible biases. Model cards help users to understand and evaluate ML models more comprehensively, such that they are not deployed in unsuited contexts, and thus to increase reproducibility. Similarly, model info sheets also provide documentation about ML models, but are specifically designed for the detection and prevention of data leakage in ML models [64]. Model info sheets are published alongside research to enable other researchers to quickly verify the validity of the data used to train ML models. They require authors to answer detailed questions about the data and corresponding train/test splits, targeting various types of data leakage [64].
An empirical study investigating twelve papers, making use of ML methods for prediction, found that a third were subject to some type of data leakage [64], and that in all those cases leakage errors could have been prevented by the use of model info sheets. Despite that, model info sheets have two main drawbacks: first, verifying the correctness of info sheets only works after reproducing the results; second, completing these sheets requires a certain level of expertise in ML. In general, model cards and model info sheets represent a promising, low-effort driver for ML reproducibility. They are especially useful in handling some of the methodological issues associated with ML models that could arise [64].